# assumeAssumption

### v1

@@ 5 Examples
Observe that $f$ fails to be regular at $(0,x_{2},x_{3})$ when $x_{2}(x_{3}-1)<0$, so the calculus rule in Corollary˜1 does not apply. Regardless, by Remark˜1, $\mathrm{lip}f(x)=0$ whenever $f(x)=0$, so it wouldn’t be of much use. We simply resort to the definition. If $(x_{1}<0\land x_{3}=1)\lor(x_{1}=0\land|x_{2}|\leq 1\land x_{3}=1)$, then $\accentset{\circ}{f}(x,r)=r^{2}$ for $r$ near 0. If $x_{1}=0\land|x_{2}|>1\land x_{3}=1$, then $\accentset{\circ}{f}(x,r)=x_{2}^{2}r^{2}>r^{2}$ near 0. Otherwise, $x_{1}>0$ and $f(x)=(x_{2}x_{1}+x_{3}-1)^{2}$, in which case
@@
and so $\lambda_{1}(\nabla^{2}f(x))=x_{2}^{2}+x_{1}^{2}+1$. By Proposition˜4, $\accentset{\circ}{f}(x,r)=(x_{2}^{2}+x_{1}^{2}+1)r^{2}/2+o(r^{2})>r^{2}$. ∎
Example 7.
@@
By ˜11, $\|f^{(4)}(x)\|=4|F^{\prime}(x)|_{2}^{4}$ when $F(x)=x_{1}x_{2}-1=0$, i.e., $\|f^{(4)}(x)\|=4(x_{1}^{2}+x_{2}^{2})^{2}$. This quantity is strictly minimized at $\pm(1,1)$ over the solution set, so one concludes by Corollary˜2. Alternatively, one notices that the flat minima are the same as that of $|x_{1}x_{2}-1|$ by using Definition˜3 and the fact that $t\in\mathbb{R}_{+}\mapsto t^{4}$ is strictly increasing. Since the new function is regular and locally Lipschitz, by Corollary˜1 it suffices to compute its Lipschitz modulus when $x_{1}x_{2}=1$, i.e., $x_{1}^{2}+x_{2}^{2}$. ∎
Example 8.
@@
Thus $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$, as claimed in the introduction. By virtue of Examples 2 and 3, it would seem futile to compute higher-order derivatives. But in this special case, the maximal eigenvectors of $f^{(2)}(x_{1},0)$ and $f^{(4)}(x_{1},0)$ align (with $(0,1)$) while $f^{(3)}(x_{1},0)=0$. Since $f^{(4)}(x_{1},0)=24x_{1}^{2}$, we deduce that $\accentset{\circ}{f}(x_{1},0,r)=r^{2}+x_{1}^{2}r^{4}+o(r^{4})$ and conclude by Definition˜3.
Example 9.
@@
Without loss of generality, assume $a_{i}\neq 0$ for all $i\in\llbracket 1,n\rrbracket$. If $I=\emptyset$, then $f(x)=(|a_{1}|x_{1}^{2}+\cdots+|a_{n}|x_{n}^{2}+1)^{2}$ and every global minimum is flat. Otherwise, let $m$ be the cardinal of $I$. The objective $f$ is invariant under the natural action of $G=\mathrm{diag}(O(m),I_{n-m})$ once we reorder the indices so that those in $I$ come first. By ˜11, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|_{2}^{2}$ when $F(x)=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}-1=0$. Since $F^{\prime}(x)=2(a_{1}x_{1},\dots,a_{n}x_{n})$, we seek to minimize $|F^{\prime}(x)|^{2}=4(a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2})$ subject to $a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$. Consider the Lagrangian $L(x,\lambda)=\sum_{i}a_{i}^{2}x_{i}^{2}+\lambda(1-\sum_{i}a_{i}x_{i}^{2})$. Since the constraint is qualified, at optimality $a_{i}^{2}x_{i}-\lambda a_{i}x_{i}=0$, namely, $a_{i}x_{i}(a_{i}-\lambda)=0$. If $\lambda\neq a_{i}$ for all $i$, then $0=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$, a contradiction. If $\lambda=a_{i}<0$ for some $i$, then $0>a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$, a contradiction. If $\lambda=a_{i}>0$ for some $i$, then $a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2}=\lambda$ and so $i\in I$. Thus, for any global minimum $\overline{x}$ of $f$ such that $\overline{x}_{i}=0$ for all $i\notin I$, $G\overline{x}$ is a strict global minimum of $\lambda_{1}(\nabla^{2}f(x))+\delta_{[f=f(\overline{x})]}$. It follows that every point in $G\overline{x}$ is flat by Corollary˜4. No other global minimum is flat by Corollary˜2. ∎
Example 10.
@@
and $\mathrm{lip}f(x)=2|x_{3}|^{2}+(|x_{1}|+|x_{2}|)^{2}$. Thus $(0,0,0)$ is the sole flat global minimum when $(a,b)=(0,0)$ by Corollary˜1. When $(a,b)\neq(0,0)$, $\mathrm{argmin}f=\{(at,bt,1/t):t\neq 0\}$. Accordingly, given $t\neq 0$, let $x_{t}=(at,bt,1/t)$ and compute
@@
This yields the expression of the flat global minima by Corollary˜1.
We now verify the claim of Proposition˜8. The objective $f$ is invariant under the natural action of the Lie group $G=\{\mathrm{diag}(t,t,1/t):t\neq 0\}$ whose Lie algebra is $\mathfrak{g}=\mathrm{span}\{\mathrm{diag}(1,1,-1)\}$. By Proposition˜6, a conserved quantity is given by $C(x)=x_{1}^{2}+x_{2}^{2}-x_{3}^{2}$. When $ab\neq 0$, the quantity
@@
By Proposition 6, a conserved quantity is given by
@@
By ˜11, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|^{2}$ when $F(x)=0$. Given $t_{1},\ldots,t_{n-1}>0$, let $x_{t}=\mathrm{diag}(t_{1}^{\upsilon_{n}},\ldots,t_{n-1}^{\upsilon_{n}},t^{-\overline{\upsilon}})\overline{x}$ where $\overline{x}$ is global minimum such that $|\overline{x}_{i}|=1$ for all $i$. We have
@@
for some constant $\omega>0$. Together with Proposition˜6, this means that ˜2 holds. Applying Corollary˜5, $\lambda_{1}(\nabla^{2}f)$ strictly decreases along the trajectories of $\nabla c$. Since $\lambda_{1}(\nabla^{2}f)+\delta_{[f=0]}$ is coercive, they must be bounded and hence globally defined by [21, Proposition 2]. Since $c$ is semi-algebraic, they must also converge to a point $\overline{x}$ where $\nabla c(\overline{x})=0$ [30], i.e., a flat global minimum. ∎
@@ Acknowledgments
I would like to thank Wenqing Ouyang for his careful reading of the manuscript and valuable feedback.

### v2

@@ 5 Examples
Observe that $f$ fails to be regular at $(0,x_{2},x_{3})$ when $x_{2}(x_{3}-1)<0$, so the calculus rule in Corollary˜3.18 does not apply. Regardless, by Remark˜3.27, $\mathrm{lip}f(x)=0$ whenever $f(x)=0$, so it wouldn’t be of much use. We simply resort to the definition. If $(x_{1}<0\land x_{3}=1)\lor(x_{1}=0\land|x_{2}|\leq 1\land x_{3}=1)$, then $\accentset{\circ}{f}(x,r)=r^{2}$ for $r$ near 0. If $x_{1}=0\land|x_{2}|>1\land x_{3}=1$, then $\accentset{\circ}{f}(x,r)=x_{2}^{2}r^{2}>r^{2}$ near 0. Otherwise, $x_{1}>0$ and $f(x)=(x_{2}x_{1}+x_{3}-1)^{2}$, in which case
@@
and so $\lambda_{1}(\nabla^{2}f(x))=x_{2}^{2}+x_{1}^{2}+1$. By Proposition˜3.17, $\accentset{\circ}{f}(x,r)=(x_{2}^{2}+x_{1}^{2}+1)r^{2}/2+o(r^{2})>r^{2}$. ∎
Example 5.2.
@@
By Fact˜3.26, $\|f^{(4)}(x)\|=4|F^{\prime}(x)|_{2}^{4}$ when $F(x)=x_{1}x_{2}-1=0$, i.e., $\|f^{(4)}(x)\|=4(x_{1}^{2}+x_{2}^{2})^{2}$. This quantity is strictly minimized at $\pm(1,1)$ over the solution set, so one concludes by Corollary˜3.19. Alternatively, one notices that the flat minima are the same as that of $|x_{1}x_{2}-1|$ by using Definition˜3.2 and the fact that $t\in\mathbb{R}_{+}\mapsto t^{4}$ is strictly increasing. Since the new function is regular and locally Lipschitz, by Corollary˜3.18 it suffices to compute its Lipschitz modulus when $x_{1}x_{2}=1$, i.e., $x_{1}^{2}+x_{2}^{2}$. ∎
Example 5.3.
@@
Thus $\lambda_{1}(\nabla^{2}f(x_{1},0))=2$, as claimed in the introduction. By virtue of Examples 3.20 and 3.21, it would seem futile to compute higher-order derivatives. But in this special case, the maximal eigenvectors of $f^{(2)}(x_{1},0)$ and $f^{(4)}(x_{1},0)$ align (with $(0,1)$) while $f^{(3)}(x_{1},0)=0$. Since $f^{(4)}(x_{1},0)=24x_{1}^{2}$, we deduce that $\accentset{\circ}{f}(x_{1},0,r)=r^{2}+x_{1}^{2}r^{4}+o(r^{4})$ and conclude by Definition˜3.2.
Example 5.4.
@@
Without loss of generality, assume $a_{i}\neq 0$ for all $i\in\llbracket 1,n\rrbracket$. If $I=\emptyset$, then $f(x)=(|a_{1}|x_{1}^{2}+\cdots+|a_{n}|x_{n}^{2}+1)^{2}$ and every global minimum is flat. Otherwise, let $m$ be the cardinal of $I$. The objective $f$ is invariant under the natural action of $G=\mathrm{diag}(O(m),I_{n-m})$ once we reorder the indices so that those in $I$ come first. By Fact˜3.26, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|_{2}^{2}$ when $F(x)=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}-1=0$. Since $F^{\prime}(x)=2(a_{1}x_{1},\dots,a_{n}x_{n})$, we seek to minimize $|F^{\prime}(x)|^{2}=4(a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2})$ subject to $a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$. Consider the Lagrangian
$L(x,\lambda)=\sum_{i=1}^{n}a_{i}^{2}x_{i}^{2}+\lambda\left(1-\sum_{i=1}^{n}a_{i}x_{i}^{2}\right).$
Since the constraint is qualified, at optimality $a_{i}^{2}x_{i}-\lambda a_{i}x_{i}=0$, namely, $a_{i}x_{i}(a_{i}-\lambda)=0$. If $\lambda\neq a_{i}$ for all $i$, then $0=a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$, a contradiction. If $\lambda=a_{i}<0$ for some $i$, then $0>a_{1}x_{1}^{2}+\dots+a_{n}x_{n}^{2}=1$, a contradiction. If $\lambda=a_{i}>0$ for some $i$, then $a_{1}^{2}x_{1}^{2}+\dots+a_{n}^{2}x_{n}^{2}=\lambda$ and so $i\in I$. Thus, for any global minimum $\overline{x}$ of $f$ such that $\overline{x}_{i}=0$ for all $i\notin I$, $G\overline{x}$ is a strict global minimum of $\lambda_{1}(\nabla^{2}f(x))+\delta_{[f=f(\overline{x})]}$. It follows that every point in $G\overline{x}$ is flat by Corollary˜3.24. No other global minimum is flat by Corollary˜3.19. ∎
Example 5.5.
@@
and $\mathrm{lip}f(x)=\sqrt{2|x_{3}|^{2}+(|x_{1}|+|x_{2}|)^{2}}$. Thus $(0,0,0)$ is the sole flat global minimum when $(a,b)=(0,0)$ by Corollary˜3.18. When $(a,b)\neq(0,0)$, $\mathrm{argmin}f=\{(at,bt,1/t):t\neq 0\}$. Accordingly, given $t\neq 0$, let $x_{t}=(at,bt,1/t)$ and compute
@@
This yields the expression of the flat global minima by Corollary˜3.18.
We now verify the claim of Proposition˜4.21. The objective $f$ is invariant under the natural action of the Lie group $G=\{\mathrm{diag}(t,t,1/t):t\neq 0\}$ whose Lie algebra is $\mathfrak{g}=\mathrm{span}\{\mathrm{diag}(1,1,-1)\}$. By Proposition˜4.11, a conserved quantity is given by $C(x)=x_{1}^{2}+x_{2}^{2}-x_{3}^{2}$. When $ab\neq 0$, the quantity
@@
By Proposition 4.11, a conserved quantity is given by
@@
By Fact˜3.26, $\lambda_{1}(\nabla^{2}f(x))=2|F^{\prime}(x)|^{2}$ when $F(x)=0$. Given $t_{1},\ldots,t_{n-1}>0$, let $x_{t}=\mathrm{diag}(t_{1}^{\upsilon_{n}},\ldots,t_{n-1}^{\upsilon_{n}},t^{-\overline{\upsilon}})\overline{x}$ where $\overline{x}$ is global minimum such that $|\overline{x}_{i}|=1$ for all $i$. We have
@@
for some constant $\omega>0$. Together with Proposition˜4.11, this means that Assumption˜4.3 holds. Applying Corollary˜4.8, $\lambda_{1}(\nabla^{2}f)$ strictly decreases along the trajectories of $\nabla c$. Since $\lambda_{1}(\nabla^{2}f)+\delta_{[f=0]}$ is coercive, they must be bounded and hence globally defined by [24, Proposition 2]. Since $c$ is semi-algebraic, they must also converge to a point $\overline{x}$ where $\nabla c(\overline{x})=0$ [35], i.e., a flat global minimum. ∎
@@ Acknowledgments
I would like to thank Wenqing Ouyang for his careful reading of the manuscript and valuable feedback. I also wish to thank Xiaopeng Li for useful criticism.